Time Frequency Brown Bag Seminar

Wednesday, March 11, 1998

12:00 noon

EQuad E415

Speaker:

Title: Wavelet-Based Video Indexing and Querying for a Smart VCR

Abstract:

We present several algorithms suitable for analysis of broadcast video. This technology could be employed to construct a ”smart VCR” that was capable of alerting the viewer to the beginning of a specific program or identifying commercials and then muting the volume on the TV. First, we show how wavelet analysis of frames of video can be used to detect transitions between shots in a video stream, thereby dividing the stream into segments. Next we describe how each segment can be inserted into a video database using an indexing scheme that involves a wavelet-based ”signature.” Finally, we show that during a subsequent broadcast of a similar video segment, the clip can be found in the database by quickly searching for the relevant signature. The method is robust against noise and typical variations in the video stream, even global changes in brightness that can fool histogrambased techniques. We compare experimentally our shot transition mechanism to a color histogram implementation, and also evaluate the effectiveness of our database searching scheme. Our algorithms are very efficient and run in real-time on a desktop computer.